OpenStudy (anonymous):

True or false Will medal ((2x-1)/(x+1))/((3x^2)/(x^2+x))=(2x-1)/(3x) And is it in reduced form

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

\[\frac{ \frac{ 2x-1 }{ x+1 } }{ \frac{ 3x^2 }{ x^2+x } }\] that

OpenStudy (danjs):

notice in the very bottom denominator, you have x^2+x which can factor to be x(x+1)

OpenStudy (danjs):

flip the bottom fraction over, and multiply it to the top fraction instead of divide

OpenStudy (danjs):

\[\frac{ 2x-1 }{ x+1 }*\frac{ x(x+1) }{ 3x^2 } = \frac{ 2x-1 }{ 3x }\]

OpenStudy (anonymous):

Yay so it's true

OpenStudy (danjs):

idk, have to continue with the left side, multiply it by

OpenStudy (danjs):

cancel out (x+1) from the top and bottom on the left

OpenStudy (danjs):

When you multiply fractions it is just \[\frac{ a }{ b }*\frac{ c }{ d } = \frac{ ac }{ bd }\]

OpenStudy (danjs):

so the left side becomes \[\frac{ x*(2x-1)(x+1) }{ 3x^2(x+1) } = \frac{ 2x-1 }{ 3x }\]

OpenStudy (danjs):

cancel common factors in the top and bottom , and see what you get

OpenStudy (danjs):

Maybe this will make it more clear \[\frac{ x*(x+1)*(2x-1) }{ 3*x*x*(x+1) }\] cancel what is common in top and bottom, then does it equal the right side of the original equation, true or false

OpenStudy (anonymous):

Yes I get 2x-1/3x

OpenStudy (danjs):

so then true

OpenStudy (anonymous):

Thank you

OpenStudy (danjs):

welcome